OpenStudy (anonymous):

solve equation (x+2)(x-3)=(x-3)^2

OpenStudy (ipwnbunnies):

Write it out: (x-2)(x-3) = (x-3)(x-3) Do you see something we can eliminate?

OpenStudy (ipwnbunnies):

(x+2)*

OpenStudy (anonymous):

Do not simplify. Develop the equation. x^2 - 3x + 2x + 6 = x^2 -6x + 9 -x + 6 = -6x + 9 5x = ...

OpenStudy (anonymous):

Expand out terms: \[x^2-x-6=x^2-6x+9\] then you have: \[5x-15=0\] \[5(x-3)=0\] \[x-3=0\] \[x=3\]